Vincent is three times as old as his brother and 6 years ago the sum of their ages was equal to Vincent's present age.How old is each?
Answer by bonster(299)

Their current ages:
Vincent=3x
Vin's Bro=x


their ages six years ago:
Vincent=3x-6
Vin's Bro=x-6
(3x-6)+(x-6)=3x
4x-12=3x
x=12



x=12-->Vincent's brother is 12 years old.
3x=36 -->Vincent is 36 years old.

Check:
(36-6)+(12-6)=36
30+6=36